OpenStudy (anonymous):

Use Gaussian Elimination to solve for A, B, C of the Partial Fraction Decomposition Equation. (x+3)/(x^2(x-1)) = A/x+B/x^2+C/(x-1) I have expanded it by multiplying by the LCD but do not know how to equate coefficients or set up equations to use Gaussian Elimination.. x+3 = A (x^2-x) +B(x-1) +Cx^2 Any help please :)))

hero (hero):

Use this to help: http://calc101.com/webMathematica/partial-fractions.jsp#topdoit

OpenStudy (anonymous):

\[x+3=A \left( x ^{2}-x \right)+B \left( x-1 \right)+C x ^{2}=\left( A+C \right)x ^{2}+\left( -A-B \right)x-B\] now equating co-efficients 0=A+C 1=-A-B 3=-B now you can find the values of A and C

OpenStudy (anonymous):

I tried using the Calc 101 website but kept getting the syntax wrong. But I will try it in the future

OpenStudy (anonymous):

@surjithayer: I found the values for A and C using your method. But how do I use Gaussian elimination from there?

OpenStudy (anonymous):